Braxton Precision Mfg. wishes to compare its inventory turnover to those of industry leaders who have turnover of about 13.5 times per year and 11.6% of their assets invested in inventory.

ADH Consolidated Net Revenue Cost of Sales Inventory Total Assets $41,250 $33,750 $2,500 $21,500 Braxton Precision Manufacturing Net Revenue Cost of Sales Inventory Total Assets $54,000 $43,000 $2,500 $33,200

a) What is Braxton's inventory turnover?

b) What is Braxton's percent of assets committed to inventory?

c) How does Braxton's performance compare to industry leaders?

Please show work.
Answer #1

a) Inventory Turnover = Cost of Sales/Average Inventory = 43000/2500 = 17.2

b) % of assets committed to inventory = (2500/33200)*100 = 7.5 %

c) Braxton Inventory Turnover is higher compared to ADH Consolidated which means it has stronger sales and is performing better.

% of assets invested in inventory is lesser compared to industry leaders which is obvious as the inventory turnover is higher.